if θ =(pi/6) and c=4,find the exact value of a and b
answer: a=2/sqrt3
b=2
how?

Answer by Alan3354(69443)

if θ =(pi/6) and c=4,find the exact value of a and b
answer: a=2/sqrt3
b=2
how?
============
If c is the hypotenuse, or the radius of a circle centered at the Origin:
a = x = c*cos(pi/6)
a = 4*(sqrt(3)/2) = 2*sqrt(3) ---- not 2/sqrt(3)
------
b = 4*sin(pi/6) = 2